### v1

# Decentralized Stochastic Subgradient Methods for Nonsmooth Nonconvex Optimization

Abstract
In this paper, we concentrate on decentralized optimization problems with nonconvex and nonsmooth objective functions, especially on the decentralized training of nonsmooth neural networks. We introduce a unified framework, named DSM, to analyze the global convergence of decentralized stochastic subgradient methods. We prove the global convergence of our proposed framework under mild conditions, by establishing that the generated sequence asymptotically approximates the trajectories of its associated differential inclusion. Furthermore, we establish that our proposed framework encompasses a wide range of existing efficient decentralized subgradient methods, including decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T), and DSGD with momentum (DSGDm). In addition, we introduce SignSGD employing the sign map to regularize the update directions in DSGDm, and show it is enclosed in our proposed framework. Consequently, our convergence results establish, for the first time, global convergence of these methods when applied to nonsmooth nonconvex objectives. Preliminary numerical experiments demonstrate that our proposed framework yields highly efficient decentralized subgradient methods with convergence guarantees in the training of nonsmooth neural networks.
Keywords: Nonsmooth optimization, decentralized stochastic subgradient method, conservative field, differential inclusion, Lyapunov function.
1 Introduction
In this paper, we consider the following decentralized optimization problem (DOP) on an undirected, connected graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$ containing $d$ agents,
$\min_{{\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{d}\in\mathbb{R}^{n}}$ $\quad\sum_{i=1}^{d}f_{i}({\bm{x}}_{i}),$ (DOP)
s. t. $\quad{\bm{x}}_{i}={\bm{x}}_{j},\quad\forall(i,j)\in\mathtt{E}.$
Here $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R},i\in[d]:=\{1,\ldots,d\}$ are nonconvex and possibly nonsmooth, and exclusively known to the agent $i$. Moreover, ${\bm{x}}_{i}$ refers to the local variables of the agent $i$. Furthermore, $\mathtt{E}$ denotes the set of edges of the graph $\mathtt{G}$ and represents communication links between agents, in the sense that all the agents communicate solely with their immediate neighbors within the graph $\mathtt{G}$.
Existing works on decentralized optimization
The decentralized optimization problem (DOP) has wide applications in data science and machine learning [12, 42], leading to increasing demands for efficient optimization methods. Motivated by stochastic gradient descent (SGD), the decentralized stochastic gradient descent (DSGD) [30, 27] is one of the most popular methods for solving (DOP). The updating scheme of DSGD can be expressed as
${\bm{x}}_{i,k+1}=\sum_{j=1}^{d}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{g}}_{i,k},$ (1.1)
where ${\bm{g}}_{i,k}$ refers to the (stochastic) gradient of $f_{i}$ at ${\bm{x}}_{i,k}$. Moreover, ${\bm{W}}\in\mathbb{R}^{d\times d}$ represents the mixing matrix corresponding to the graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$. That is, ${\bm{W}}$ is symmetric, doubly stochastic, and ${\bm{W}}(i,j)=0$ whenever $(i,j)\notin\mathtt{E}$ (see Section 2.2 for details). By denoting the collection of local variables at $k$-th iteration as ${\bm{X}}_{k}\in\mathbb{R}^{n\times d}$, where
${\bm{X}}_{k}:=\left[{\bm{x}}_{1,k},{\bm{x}}_{2,k},\ldots,{\bm{x}}_{d,k}\right],$ (1.2)
and with ${\bm{G}}_{k}=[{\bm{g}}_{1,k},{\bm{g}}_{2,k},\ldots,{\bm{g}}_{d,k}]\in\mathbb{R}^{n\times d}$, we can reformulate the update equation (1.1) into the following compact form,
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{G}}_{k}.$ (1.3)
Based on the update scheme of DSGD, a wide range of efficient decentralized methods have been developed for solving (DOP) under the assumption on differentiability of its objective functions. For example, [49, 19] proposes the DSGD with momentum (DSGDm) by integrating the heavy-ball momentum [32] into the DSGD. Moreover, some recent works introduce auxiliary variables within DSGD to track the exact gradient of the objective function in each agent, such as the DSGD with gradient-tracking technique (DSGD-T) [48], EXTRA [37], etc. To provide theoretical guarantees for these decentralized methods with continuously differentiable objective functions, [27] provides the first theoretical analysis for the convergence rate of DSGD. Moreover, [52, 34, 23, 1, 50] further explore the convergence properties of the DSGD, especially regarding network topology and data heterogeneity. Additionally, [49, 51, 28] analyze the global convergence of the different versions of DSGDm they proposed, while [53, 47, 33, 39] show the convergence properties of DSGD-T for continuously differentiable objectives. Furthermore, some existing works develop decentralized proximal gradient methods for (DOP), under the assumption that $f_{i}$ is composed of a differentiable term $g_{i}$ and a convex regularization term $r_{i}$ for all $i\in[d]$. In this context, the objective function is weakly convex and thus Clarke regular. For example, [52, 38] introduces the proximal decentralized gradient method and establishes its global convergence properties based on the update scheme of DSGD. Moreover, [17] introduces a decentralized proximal method named NEXT based on similar gradient-tracking techniques as DSGD-T. Furthermore, [35, 24] improve the convergence analysis for DSGD-T and prove its convergence for weakly convex objectives. Although achieving nice theoretical results, all the aforementioned works require the Clarke regularity [14] of the objective functions in (DOP). However, as discussed in [8], the nonsmooth activation functions, including ReLU and leaky ReLU, are popular building blocks for model neural networks. More importantly, in the training of neural networks, the widely employed ReLU activation function leads to the nonsmoothness of the loss function. As a result, in the decentralized training of various popular neural networks, the objective functions $\{f_{i}:1\leq i\leq d\}$ in (DOP) are usually nonsmooth and lacking Clarke regularity. When applied to the decentralized training of these nonsmooth neural networks, all the aforementioned works are not capable of providing convergence guarantees for their analyzed decentralized methods. How to establish convergence properties for decentralized stochastic subgradient methods under nonconvex and nonsmooth settings largely remains unexplored.
Stochastic subgradient methods for nonsmooth nonconvex optimization
In the training of nonsmooth neural networks, one major challenge comes from how to compute the (stochastic) subgradients of the loss function. In a wide range of deep learning packages, the subgradients of the loss function are computed by the automatic differentiation (AD) algorithms. These algorithms, utilizing the chain rule, recursively construct generalized subgradients by composing the Jacobians of each block. However, since the chain rule fails for Clarke subdifferential [14], the outputs of the AD algorithm may not be contained in the Clarke subdifferential of the loss function. To characterize the outputs of AD algorithms in differentiating nonsmooth neural networks, [15, 8] introduce the class of path-differentiable functions, which is general enough to cover a wide range of objective functions in real-world applications, including semi-algebraic functions, semi-analytic functions, and functions whose graphs are definable in some o-minimal structures [15]. For path-differentiable functions, [8] introduces the concept of the conservative field as a generalization of their Clarke subdifferential, and shows that the chain rule holds for conservative fields. As a result, although the results of AD algorithms may not be constrained in the Clarke subdifferential of the loss function of a nonsmooth neural network, these results are contained in a specific conservative field of that loss function. Therefore, we employ the concept of the conservative field as a generalization of the Clarke subdifferential to characterize the outputs of AD algorithms throughout our paper. To analyze the convergence properties of stochastic subgradient methods, the ordinary differential equation (ODE) approaches are developed [2, 10, 18, 15], which describe the convergence of the iterates in a subgradient method through the asymptotic behavior of the dynamics of corresponding differential inclusion (i.e., a nonsmooth generalization of ODE). Then based on the concepts of path-differentiable functions and conservative field, a wide range of stochastic subgradient methods are developed for nonsmooth nonconvex optimization problems on a single agent (i.e., (DOP) with $d=1$), while enjoying convergence guarantees from ODE approaches. [15, 7, 8] study the global convergence of SGD method. Moreover, [36, 26, 46] introduce the SGD with heavy-ball momentum and establish its global convergence. In addition, [11] proposes the inertial Newton method (INNA) and shows its global convergence on training nonsmooth neural networks. Furthermore, [31] proves the global convergence of the AdaGrad-Norm, and [45] provides convergence guarantees for a family of ADAM methods. In addition, [21, 22] discuss stochastic subgradient methods for optimization problems over manifolds. Significant progress has been made in the study of stochastic subgradient descent methods within a single agent. This motivates us to explore decentralized stochastic subgradient methods under nonconvex and nonsmooth settings based on the ODE approaches mentioned above. Although [41] establishes the global convergence of DSGD on nonsmooth functions utilizing the ODE approach, their analysis heavily relies on the condition $\partial(\sum_{i\in[d]}f_{i})=\sum_{i\in[d]}\partial f_{i}$. This assumption presumes the validity of the chain rule for $\{f_{i}\}$ in (DOP). As the chain rule does not hold for the Clarke subdifferential of general nonsmooth functions, hence limiting their analysis to Clarke regular functions [14]. How to establish convergence properties for decentralized subgradient methods in minimizing nonsmooth path-differentiable functions remains a compelling open question for further investigations.
Contributions
The main contributions of our paper are summarized as follows:
  • A novel framework for decentralized stochastic subgradient methods We introduce set-valued mappings $\{\mathcal{H}_{i}:i\in[d]\}$, where $\mathcal{H}_{i}:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ characterizes the update schemes within the $i$-th agent. Then we first introduce the following framework based on the DSGD in (1.3),
    ${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}).$ (DSM)
    Here for any $i\in[d]$, ${\bm{Z}}_{i,k}\in\mathbb{R}^{m}$ denotes the local copies of the variables at the $i$-th agent, and ${\bm{Z}}_{k}=[{\bm{z}}_{1,k},\ldots,{\bm{z}}_{d,k}]\in\mathbb{R}^{m\times d}$, i.e., the nested local variables. Moreover, ${\bm{H}}_{k}\in\mathbb{R}^{m\times d}$ refers to the noiseless nested update directions of all the agents, in the sense that ${\bm{H}}_{k}{\bm{e}}$ is an approximated evaluation of $\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}}_{i,k})$ with ${\bm{e}}:=[\frac{1}{d},\ldots,\frac{1}{d}]^{\top}\in\mathbb{R}^{d}$. We show that such a condition is general enough to enclose a wide range of decentralized subgradient methods. Furthermore, $\Xi_{k+1}\in\mathbb{R}^{m\times d}$ refers to the noises, and $\eta_{k}>0$ denotes the step-size. In addition, ${\bm{W}}\in\mathbb{R}^{d\times d}$ refers to the mixing matrix that characterizes the local consensus of all the agents in each iteration. We prove that under mild conditions, $\lim_{k\to+\infty}\left\|{\bm{z}}_{i,k}-{\bm{z}}_{j,k}\right\|=0$ holds for any $i,j\in[d]$, and the sequence $\{{\bm{z}}_{i,k}\}$ asymptotically approximates the trajectories of the differential inclusion
    $\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}})\right).$ (1.4)
    When the differential inclusion (1.4) admits Lyapunov function $\Psi$ and stable set $\mathcal{A}$, we prove that any cluster point of $\{{\bm{Z}}_{i,k}\}$ lies in $\mathcal{A}$ under mild conditions based on the existing ODE approaches.
  • Decentralized stochastic subgradient methods with convergence guarantees We demonstrate that a wide range of decentralized stochastic subgradient methods fit into our proposed framework, including DSGD, DSGDm, and DSGD-T. Furthermore, we also propose a decentralized variant of SignSGD denoted as DSignSGD which can be enclosed by (DSM). More importantly, based on our proposed framework, we provide convergence guarantees for these decentralized subgradient methods. As there is no existing work on the convergence properties of DSGDm, DSGD-T, and DSignSGD in nonsmooth nonconvex optimization, our results fill the gap between their practical implementations and theoretical guarantees. Preliminary numerical experiments show the efficiency of analyzed methods and exhibit the superiority of our proposed method DSignSGD in comparison with DSGD and DSGDm, hence demonstrating the flexibility and great potential of our framework.
Organizations
The rest of this paper is organized as follows. In section 2, we provide an overview of the notations that appeared throughout the paper and present the preliminary concepts and properties of mixing matrix, conservative field, and differential inclusion. In section 3, we revisit our proposed framework (DSM) for decentralized stochastic subgradient methods, and establish the global convergence by linking it with the dynamics of its corresponding differential inclusion. In Section 4 we demonstrate how several decentralized stochastic subgradient methods can be unified within our framework, and develop the convergence properties for these methods. We present numerical experiments of analyzed decentralized subgradient methods in decentralized training nonsmooth neural networks in Section 5. In the last section, We draw conclusions and discuss some possible future research directions.
2 Preliminaries
2.1 Notations
[…]
$f({\bm{x}}):=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ (2.1)
and $F({\bm{X}}):=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}}_{i})$ with ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{d}]$. For any positive sequence $\left\{\eta_{k}\right\}$, let $\lambda(0):=0,\lambda(i):=\sum_{k=0}^{i-1}\eta_{k}$, and $\Lambda(t):=\sup\left\{k\in\mathbb{N}:t\geq\lambda(k)\right\}.$ More explicitly, $\Lambda(t)=p$, if $\lambda(p)\leq t<\lambda(p+1)$. Furthermore, we define the set-valued mapping sign : $\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as follows:
$(\operatorname{sign}({\bm{x}}))_{i}=\begin{cases}\{-1\}&{\bm{x}}_{i}<0\\ {[-1,1]}&{\bm{x}}_{i}=0\\ \{1\}&{\bm{x}}_{i}>0\end{cases}$
Then it is easy to verify that $\operatorname{sign}({\bm{x}})$ is the Clarke subdifferential of the nonsmooth functions ${\bm{x}}\mapsto\|{\bm{x}}\|_{1}$. In addition, we denote $\mathcal{N}_{i}$ as the set of neighbor agents of agent $i$ together with $\{i\}$, $\mathbb{R}_{+}$ as the set of all nonnegative real numbers. For any $N>0,[N]:=\{1,\ldots,N\}$. And let ${\bm{e}}:=\frac{1}{d}{\bm{1}}_{d}$, where ${\bm{1}}_{d}$ represents all-ones vector. Furthermore, we denote $(\Omega,\mathcal{F},\mathbb{P})$ as the probability space. We say $\left\{\mathcal{F}_{k}\right\}_{k\in\mathbb{N}}$ is a filtration if $\left\{\mathcal{F}_{k}\right\}$ is a collection of $\sigma$-algebras that satisfies $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{\infty}\subseteq\mathcal{F}$. A stochastic series $\{\xi_{k}\}$ is called a martingale with respect to a filtration $\{\mathcal{F}_{k}\}$, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=\xi_{k}$, $\forall k\in\mathbb{N}$. Moreover, A stochastic series $\{\xi_{k}\}$ is referred to as a martingale difference sequence, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=0$ holds for $\forall k\in\mathbb{N}$.
2.2 Mixing matrix
In this subsection, we introduce the concept of mixing matrix ${\bm{W}}=[{\bm{W}}(i,j)]\in\mathbb{R}^{d\times d}$, which determines the structure of the communication network and plays an important role in averaging the local information of neighbor agents. In general, we assume the mixing matrix ${\bm{W}}$ satisfies the following properties, which are standard in the literature.
Definition 2.1.
Given a connected graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$, we say ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix of $\mathtt{G}$, if it satisfies the following conditions.
  1. 1.
    ${\bm{W}}$ is symmetric.
  2. 2.
    ${\bm{W}}$ is doubly stochastic, namely, ${\bm{W}}$ is nonnegative and ${\bm{W}}{\bm{1}}_{d}={\bm{W}}^{\top}{\bm{1}}_{d}={\bm{1}}_{d}$.
  3. 3.
    ${\bm{W}}(i,j)=0$, if and only if $i\neq j$ and $(i,j)\notin\mathtt{E}$.
With a given graph $\mathtt{G}$, there are various approaches to choose its corresponding mixing matrix, such as Laplacian-based constant edge weight matrix [43] and Metropolis weight matrix [44]. We refer interested readers to [29, 37] for more details about how to choose the mixing matrix. Based on the Definition 2.1, we present the following proposition to illustrate the fundamental properties of the mixing matrix. The proof of Proposition 2.2 follows from direct calculations, hence is omitted for simplicity.
Proposition 2.2.
For any mixing matrix ${\bm{W}}\in\mathbb{R}^{d\times d}$ that corresponds to a connected graph $\mathtt{G}$, all the eigenvalues of ${\bm{W}}$ lie in $(-1,1]$, and ${\bm{W}}$ has a single eigenvalue at $1$ that admits ${\bm{e}}$ as its eigenvector.
2.3 Nonsmooth analysis and conservative field
In this subsection, we present a brief introduction to some basic concepts in nonsmooth analysis and the conservative field [8]. A set-valued mapping $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is a mapping from $\mathbb{R}^{n}$ to set of subsets of $\mathbb{R}^{m}$. The graph of $\mathcal{G}$ is given by
$\operatorname{graph}\mathcal{G}=\{({\bm{x}},{\bm{z}})\mid{\bm{x}}\in\mathbb{R}^{n},{\bm{z}}\in\mathcal{G}({\bm{x}})\}.$
The set-valued mapping is said to be graph-closed (or has closed graph), if $\operatorname{graph}\mathcal{G}$ is a closed subset of $\mathbb{R}^{n}\times\mathbb{R}^{m}$. $\mathcal{G}$ is said to be locally bounded, if for any ${\bm{x}}\in\mathbb{R}^{n}$, there exists a neighborhood $U_{{\bm{x}}}$ of ${\bm{x}}$ such that
$\sup\limits_{{\bm{z}}\in\mathcal{G}({\bm{y}}),{\bm{y}}\in U_{{\bm{x}}}}\|{\bm{z}}\|<+\infty.$
In addition, $\mathcal{G}$ is convex-valued, if $\mathcal{G}({\bm{x}})$ is a convex subset of $\mathbb{R}^{n}$ for any ${\bm{x}}\in\mathbb{R}^{n}$. Moreover, A curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is absolutely continuous if $\gamma$ is differentiable almost everywhere, and $\gamma(t)=\gamma(0)+\int_{0}^{t}\gamma^{\prime}(u)\mathrm{d}u$ holds for all $t\in\mathbb{R}_{+}$. For any locally Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$, Rademacher’s theorem states that a locally Lipschitz continuous function $f$ is differentiable almost everywhere. We denote the subset in $\mathbb{R}^{n}$ on which $f$ is differentiable by $\Omega_{f}$. According to [13, Theorem 2.5.1], we have
$\partial{f({\bm{x}})}=\operatorname{conv}\left\{\lim_{i\to+\infty}\nabla f\left({\bm{x}}_{i}\right):{\bm{x}}_{i}\rightarrow{\bm{x}},~{}{\bm{x}}_{i}\in\Omega_{f}\right\}.$
Then it is easy to verify that $\partial f$ is a convex-valued graph-closed locally bounded set-valued mapping from $\mathbb{R}^{n}$ to the subsets of $\mathbb{R}^{n}$. Now we present a brief introduction to the conservative field.
Definition 2.3 (Conservative field).
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a nonempty set-valued mapping. We call $\mathcal{G}$ a conservative field if it is compact-valued and graph closed, and for any absolutely continuous loop $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$ satisfying $\gamma(0)=\gamma(1)$, it holds that
$\int_{0}^{1}\max_{{\bm{v}}\in\mathcal{G}(\gamma(t))}\langle\dot{\gamma}(t),{\bm{v}}\rangle\mathrm{d}t=0.$ (2.2)
We now introduce the definition of the potential function corresponding to the conservative field.
Definition 2.4.
Let $\mathcal{G}$ be a conservative field, we say a function $f$ is a potential for $\mathcal{G}$, provided that $f$ can be formulated as
$f({\bm{x}})={f({\bm{x}}_{0})}+\int_{0}^{t}\langle\dot{\gamma}(s),\mathcal{G}(\gamma(s))\rangle\mathrm{d}s,$
for any absolutely continuous curve $\gamma$ with $\gamma(0)={\bm{x}}_{0}$ and $\gamma(t)={\bm{x}}$. We can also say that $\mathcal{G}$ is a conservative field for $f$. Such functions $f$ are called path-differentiable functions.
As illustrated in [8, Corollary 1], for any given path-differentiable function $f$, $\partial f$ is a conservative field for $f$. Therefore, the concept of the conservative field can be regarded as an extension of the Clarke subdifferential. At the last of this part, we provide some explanations about the path-differentiable functions. In fact, the class of path-differentiable functions is general enough to cover the most objective functions in real-world problems. The well-known Clarke regular functions mentioned in [13] and semi-algebraic functions mentioned in [54] are all path-differentiable functions. In [16], the authors discuss an important function class of path-differentiable functions, definable functions, which are referred to as the functions whose graphs are definable in an o-minimal structure. Actually, most activation functions and loss functions in deep neural networks are definable functions, including sigmoid, softplus, ReLU, leaky ReLU, hinge loss, etc. Furthermore, it is important to highlight that definability remains preserved under finite summation and composition [15]. Consequently, for any neural network constructed by definable blocks, its objective function is definable, thus it is a path-differentiable function. Additionally, it is worth noting that the Clarke subdifferential of definable functions is itself definable [6]. Hence, in the case of a neural network constructed by definable blocks, the conservative field corresponding to the automatic differentiation algorithms is also definable. The following proposition demonstrates that the definability of both $f$ and its conservative field $\mathcal{G}$ leads to the nonsmooth Morse–Sard property [5].
Proposition 2.5 (Theorem 5 in [6]).
Let $f$ be a path-differentiable function that admits $\mathcal{D}_{f}$ as its conservative field. Suppose both $f$ and $\mathcal{D}_{f}$ are definable over $\mathbb{R}^{n}$, then the set $\left\{{f({\bm{x}})}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\right\}$ is finite.
Finally, based on the concept of the conservative field, we present the definition of the critical points for the optimization problem (DOP),
Definition 2.6.
Let $f$ in (2.1) be a path-differentiable function that admits $\mathcal{D}_{f}$ as its convex-valued conservative field, then we say ${\bm{X}}\in\mathbb{R}^{n\times d}$ is a $\mathcal{D}_{f}$-critical point of (DOP), if ${\bm{X}}={\bm{X}}{\bm{e}}{\bm{1}}_{d}^{\top}$ and ${\bm{0}}\in\mathcal{D}_{f}({\bm{X}}{\bm{e}})$.
2.4 Differential inclusion and stochastic approximation
In this subsection, we introduce fundamental concepts and theories associated with differential inclusion, which is crucial for establishing the convergence properties of stochastic approximation. We begin with the concept of differential inclusion.
Definition 2.7.
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a set-valued mapping. An absolutely continuous curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is called a solution (or trajectory) to the differential inclusion
$\frac{d{\bm{x}}}{dt}\in\mathcal{G}({\bm{x}}),$ (2.3)
(or a trajectory of $\mathcal{G}$) with initial point ${\bm{x}}$, provided that $\gamma(0)={\bm{x}}$ and $\gamma^{\prime}(t)\in\mathcal{G}(\gamma(t))$ holds for almost every $t\in\mathbb{R}_{+}$.
We now present the definitions of the Lyapunov function and perturbed solution.
Definition 2.8 (Lyapunov function).
Let $\mathcal{B}\subset\mathbb{R}^{n}$ be a closed set. A continuous function ϕ:nnormal-:italic-$\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is referred to as a Lyapunov function for the differential inclusion (2.3) with the stable set $\mathcal{B}$, if for any solution $\gamma$ to (2.3) and any $t>0$, it holds that
$\phi(\gamma(t))\leq\phi(\gamma(0)).$
Moreover, whenever $\gamma(0)\notin\mathcal{B}$, it holds for any $t>0$ that
$\phi(\gamma(t))<\phi(\gamma(0)).$
Definition 2.9.
For any set-valued mapping $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ and any $\delta\geq 0$, we denote $\mathcal{G}^{\delta}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as
$\mathcal{G}^{\delta}({\bm{x}})=\cup_{{\bm{y}}\in\mathbb{B}_{\delta}({\bm{x}})}(\mathcal{G}({\bm{y}})+\mathbb{B}(\delta,{\bm{0}})).$
Definition 2.10 (Perturbed solution).
An absolutely continuous curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is said to be a perturbed solution of differential inclusion (2.3), if the following two conditions hold:
  • There exists a locally integrable function $w:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$, such that
[…]
    holds for any $T>0$.
  • For above locally integrable $w$, there exists $\delta:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}\delta(t)=0$ and
    $\dot{\gamma}(t)-w(t)\in\mathcal{G}^{\delta(t)}(\gamma(t)).$
Then, we consider a discrete iterative sequence corresponding to the differential inclusion (2.3),
${\bm{x}}_{k+1}={\bm{x}}_{k}+\eta_{k}({\bm{y}}_{k}+\xi_{k+1}),$ (2.4)
where $\{\eta_{k}\}$ is a diminishing positive sequence of step-sizes, ${\bm{y}}_{k}$ is an approximate estimate of $\mathcal{G}$ at ${\bm{x}}_{k}$, and $\xi_{k+1}$ is a random noise when evaluating ${\bm{y}}_{k}$. A continuous time interpolated process $u:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ can be induced by (2.4),
$u(\lambda(k)+s)={\bm{x}}_{k}+\frac{{\bm{x}}_{k+1}-{\bm{x}}_{k}}{\eta_{k}}s,\quad s\in[0,\eta_{k}).$ (2.5)
Here, $\lambda(0):=0$ and $\lambda(n):=\sum_{i=0}^{n-1}\eta_{i},n\geq 1$. The following lemma shows a sufficient condition under which the interpolated process is a perturbed solution of (2.3).
Lemma 2.11 (Lemma 2.20 in [46]).
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded convex-valued graph-closed set-valued mapping. Suppose the following conditions hold in (2.4):
  1. 1.
    For any $T>0$, it holds that
    $\lim\limits_{m\rightarrow+\infty}\sup\limits_{m\leq s\leq\Lambda(\lambda(m)+T)}\left\|\sum_{k=m}^{s}\eta_{k}\xi_{k+1}\right\|=0.$
  2. 2.
    There exists a diminishing nonnegative sequence $\{\delta_{k}\}$ such that $d_{k}\in\mathcal{G}^{\delta_{k}}({x_{k}})$ holds for all $k\geq 0$.
  3. 3.
    It holds that $\sup_{k\geq 0}\left\|{\bm{x}}_{k}\right\|<+\infty$, and $\sup_{k\geq 0}\left\|{\bm{y}}_{k}\right\|<+\infty$.
Then the interpolated process $u(t)$ of $\{{\bm{x}}_{k}\}$ is a perturbed solution of (2.3).
The following theorem summarizes the results in [2], which illustrates the convergence of $\{{\bm{x}}_{k}\}$ generated by (2.4). Therefore, we omit the proof of Theorem 2.12 for simplicity.
Theorem 2.12.
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded set-valued mapping that is nonempty compact convex valued with closed graph. For any sequence $\{{\bm{x}}_{k}\}$, suppose there exists a continuous function Ψ:nnormal-:normal-$\Psi:\mathbb{R}^{n}\to\mathbb{R}$ and a closed subset $\mathcal{A}$ of $\mathbb{R}^{n}$ such that
  1. 1.
    $\Psi$ is bounded from below, and the set $\{\Psi({\bm{x}})\,:\,{\bm{x}}\in\mathcal{A}\}$ has empty interior in $\mathbb{R}$.
  2. 2.
    $\Psi$ is a Lyapunov function for the differential inclusion (2.3) that admits $\mathcal{A}$ as its stable set.
  3. 3.
    The interpolated process of $\{{\bm{x}}_{k}\}$ is a perturbed solution of (2.3).
Then any cluster point of $\{{\bm{x}}_{k}\}$ lies in $\mathcal{A}$, and the sequence $\{\Psi({\bm{x}}_{k})\}$ converges.
Similar results under slightly different conditions can be found in [10, 15, 18], while some recent works [4, 9] focus on analyzing the convergence of (2.4) under relaxed conditions. Interested readers could refer to those works for details.
3 A General Framework for Nonsmooth Decentralized Subgradient Methods
In this section, we aim to establish the global convergence of the framework (DSM). Section 3.1 exhibits the fundamental assumptions in our analysis, and in Section 3.2 we present the convergence properties of the framework (DSM).
3.1 Basic assumptions
In this subsection, we introduce the basic assumptions of our theoretical analysis. To begin with, we make some mild assumptions on our framework (DSM).
Assumption 3.1.
  1. 1.
    ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix.
  2. 2.
    The sequence $\{{\bm{Z}}_{k}\}$ and $\{{\bm{H}}_{k}\}$ are bounded.
  3. 3.
    The sequences of step-sizes $\{\eta_{k}\}$ is positive and diminishing, which satisfies $\sum_{k=0}^{+\infty}\eta_{k}=+\infty$. Moreover, for any $T>0$, it holds that
    $\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$ (3.1)
  4. 4.
    There exists a positive diminishing sequence $\{\delta_{k}\}$ and a locally bounded graph-closed set-valued mapping $\mathcal{H}_{i}$, such that for each $i\in[d]$, it holds that
    ${\bm{H}}_{k}{\bm{e}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right),\quad\forall k\in\mathbb{N}.$ (3.2)
Here we present some comments on Assumption 3.1. Assumption 3.1(2) assumes the uniform boundedness of the sequence $\{{\bm{Z}}_{k}\}$, which is commonly employed in existing works [8, 15, 46, 45]. Assumption 3.1(3) illustrates that the noises in the framework (DSM) can be controlled by the step-sizes, which is identical to the condition on the noises in [2]. Furthermore, Assumption 3.1(4) illustrates that the averaged update directions of all the agents $\{{\bm{H}}_{k}{\bm{e}}\}$ approximate the evaluation of $\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}}_{i,k})$. For conciseness, we define the set-valued mapping $\mathcal{H}:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ as
$\mathcal{H}({\bm{z}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}})\right).$ (3.3)
Then we present the following assumptions on the differential inclusion
$\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\mathcal{H}({\bm{z}}),$ (3.4)
which corresponds to the discrete update scheme in framework (DSM).
Assumption 3.2.
  1. 1.
    There exists a locally Lipschitz continuous Lyapunov function $\Psi:\mathbb{R}^{m}\to\mathbb{R}$ for the differential inclusion (3.4) with stable set $\mathcal{A}$.
  2. 2.
    The set $\{\Psi({\bm{z}}):{\bm{z}}\in\mathcal{A}\}$ has empty interior in $\mathbb{R}$.
Assumption 3.2(1) clarifies the Lyapunov properties of a differential inclusion with respect to $-\mathcal{H}$, which is standard in the literature and can be satisfied by most appropriately selected set-valued mappings. Assumption 3.2(2) is the Weak Sard’s theorem, which stipulates that the set of values of $\Psi$ at points outside the stable set is dense in $\mathbb{R}$. As demonstrated in demonstrated in [15, 11], Assumption 3.2(2) is mild in practice.
3.2 Global convergence
In this subsection, we establish the convergence properties of our proposed framework (DSM), under Assumption 3.1 and Assumption 3.2. Towards this end, we first conduct consensus analysis of $\{{\bm{Z}}_{k}\}$. Denote
${\bm{P}}={\bm{e}}{\bm{1}}_{d}^{\top},\quad{\bm{P}}_{\perp}={\bm{I}}_{d}-{\bm{e}}{\bm{1}}_{d}^{\top},$
then any matrix ${\bm{Z}}\in\mathbb{R}^{m\times d}$ has the following orthogonal decomposition,
${\bm{Z}}={\bm{Z}}{\bm{P}}+{\bm{Z}}{\bm{P}}_{\perp}.$ (3.5)
Define the averaging vector $\hat{{\bm{Z}}}:={\bm{Z}}{\bm{e}}$, and the disagreement matrix ${\bm{Z}}_{\perp}:={\bm{Z}}{\bm{P}}_{\perp}$. Then (3.5) can be rephrased as
${\bm{Z}}=\hat{{\bm{Z}}}{\bm{1}}_{d}^{\top}+{\bm{Z}}_{\perp}$ (3.6)
We begin our proof with the following preliminary lemma from [40, Lemma 3.1].
Lemma 3.3.
Let $\{\gamma_{k}\}$ be a sequence of real numbers. If $10<\alpha<$ 1 and $\lim_{k\rightarrow+\infty}\gamma_{k}=\gamma$, then it holds that
$\lim_{k\rightarrow+\infty}\sum_{\ell=0}^{k}\alpha^{k-\ell}\gamma_{\ell}=\frac{\gamma}{1-\alpha}.$
Next, we present the following lemma to show that $\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{e}}\right\|$ converges to 00 as $k$ goes to infinity.
Lemma 3.4.
Suppose Assumption 3.1 holds, then for any $\{{\bm{Z}}_{k}\}$ generated by the framework (DSM), we have $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$. Moreover, it holds that
$\lim_{k\to+\infty}\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{e}}\right\|=0.$ (3.7)
Proof.
From Assumption 3.1(2) and the fact that $\lim_{k\to+\infty}\eta_{k}=0$ in Assumption 3.1(3), it holds that $\limsup_{k\to+\infty}\eta_{k}\left\|{\bm{H}}_{k}\right\|=0$. Furthermore, Assumption 3.1(3) illustrates that
$\limsup_{i\to+\infty}\left\|\eta_{i}\Xi_{i+1}\right\|\leq\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$ (3.8)
Therefore, we can conclude that
$\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0.$ (3.9)
Moreover, from the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM) and Definition 2.1, it holds that
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}.$ (3.10)
Together with (3.9), it holds that
$\lim_{k\to+\infty}\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{e}}\right\|=\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}\right\|\leq\lim_{k\to+\infty}\frac{1}{\sqrt{d}}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0.$ (3.11)
This completes the proof. ∎
Next, we denote ${\bm{Z}}_{\perp,k}:={\bm{Z}}_{k}{\bm{P}}_{\perp}$ as the disagreement matrix at $k$-th iteration. Intuitively, ${\bm{Z}}_{\perp,k}$ measures the dissimilarity of the local variables of all the agents at $k$-th iteration. In the following proposition, we prove that $\lim_{k\to+\infty}{\bm{Z}}_{\perp,k}=0$, i.e., the dissimilarity of local variables of all the agents converges to 00.
Proposition 3.5.
Suppose Assumption 3.1 holds, then for any $\{{\bm{Z}}_{k}\}$ generated by the framework (DSM), it holds that
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{\perp,k}\right\|=0.$ (3.12)
Proof.
From Definition 2.1, we can conclude that there exists $\alpha\in(0,1]$ such that $\left\|{\bm{W}}({\bm{I}}_{d}-{\bm{P}})\right\|=1-\alpha$. Then for any $k\geq 0$, it follows from the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM) that
$\left\|{\bm{Z}}_{\perp,k+1}\right\|=\left\|[{\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})]({\bm{I}}_{d}-{\bm{P}})\right\|$ (3.13)
[…]
Here the second inequality follows from the fact that $({\bm{I}}_{d}-{\bm{P}}){\bm{W}}({\bm{I}}_{d}-{\bm{P}})={\bm{W}}({\bm{I}}_{d}-{\bm{P}})$. Since Lemma 3.4 illustrates that $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$, then it can be readily deduced by Lemma 3.3 that
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{\perp,k}\right\|=0.$ (3.14)
This completes the proof. ∎
As illustrated in Proposition 3.5, any sequence $\{{\bm{Z}}_{k}\}$ from (DSM) satisfies
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{P}}\right\|=\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{e}}{\bm{1}}_{d}^{\top}\right\|=0.$ (3.15)
As a result, we can conclude that the cluster points of the sequence $\{{\bm{Z}}_{k}\}$ are exactly the same as those of the sequence $\{{\bm{Z}}_{k}{\bm{e}}{\bm{1}}_{d}^{\top}\}$. To analyze the convergence properties of $\{{\bm{Z}}_{k}{\bm{e}}\}$, we present the following proposition to illustrate the update scheme of $\{{\bm{Z}}_{k}{\bm{e}}\}$.
Proposition 3.6.
Suppose Assumption 3.1 holds, then there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that for any sequence $\{{\bm{Z}}_{k}\}$ generated from the framework (DSM), it holds for all $k\geq 0$ that
${\bm{Z}}_{k+1}{\bm{e}}\in{\bm{Z}}_{k}{\bm{e}}-\eta_{k}\left(\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}})+\Xi_{k+1}{\bm{e}}\right).$ (3.16)
Proof.
From the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM), we can conclude that
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{W}}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}={\bm{Z}}_{k}{\bm{e}}-\eta_{k}{\bm{H}}_{k}{\bm{e}}-\eta_{k}\Xi_{k+1}{\bm{e}}.$ (3.17)
Let $\delta^{\star}_{k}=\left\|{\bm{Z}}_{k}({\bm{I}}_{d}-{\bm{P}})\right\|$, from the definition of ${\bm{P}}$, it holds that $\left\|{\bm{z}}_{i,k}-{\bm{Z}}_{k}{\bm{e}}\right\|\leq\left\|{\bm{Z}}_{k}({\bm{I}}_{d}-{\bm{P}})\right\|=\delta^{\star}_{k}$. Together with Assumption 3.1(3), we can conclude that
${\bm{H}}_{k}{\bm{e}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right)\subseteq\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}+\delta^{\star}_{k}}({\bm{Z}}_{k}{\bm{e}})\right).$ (3.18)
Now we claim that there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that
$\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}+\delta^{\star}_{k}}({\bm{Z}}_{k}{\bm{e}})\right)\subseteq\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}}).$ (3.19)
Otherwise, there exists $\delta_{\varepsilon}>0$ and a subsequence $\{k_{j}\}\subset\mathbb{N}_{+}$ such that
$\sup\left\{\mathrm{dist}\left({\bm{y}},\mathcal{H}^{\delta_{\varepsilon}}({\bm{Z}}_{k_{j}}{\bm{e}})\right):{\bm{y}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}({\bm{Z}}_{k_{j}}{\bm{e}})\right)\right\}>0.$ (3.20)
From the uniform boundedness of $\{{\bm{Z}}_{k}\}$ in Assumption 3.1(1), without loss of generality, we can assume that $\{{\bm{Z}}_{k_{j}}\}$ converges to $\tilde{{\bm{Z}}}\in\mathbb{R}^{n\times d}$. Then from the graph closedness and local boundedness of $\mathcal{H}_{i}$, we have
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\mathcal{H}_{i}(\tilde{{\bm{Z}}}{\bm{e}})\right):{\bm{y}}\in\mathcal{H}_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}({\bm{Z}}_{k_{j}}{\bm{e}})\right\}=0.$ (3.21)
Then based on Jensen’s inequality, it holds that
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\mathcal{H}(\tilde{{\bm{Z}}}{\bm{e}})\right):{\bm{y}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}({\bm{Z}}_{k_{j}}{\bm{e}})\right)\right\}=0.$ (3.22)
which contradicts (3.20). As a result, there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that (3.19) holds, hence
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{W}}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}\in{\bm{Z}}_{k}{\bm{e}}-\eta_{k}\left(\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}})+\Xi_{k+1}{\bm{e}}\right).$ (3.23)
This completes the proof. ∎
Proposition 3.7.
Suppose Assumption 3.1 holds. Then the interpolated process of the sequence $\{{\bm{Z}}_{k}{\bm{e}}\}$ is a perturbed solution to the differential inclusion (3.4).
Proof.
Firstly, from Proposition 3.6 we can conclude that there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that the sequence $\{{\bm{Z}}_{k}{\bm{e}}\}$ follows the update scheme
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{W}}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}\in{\bm{Z}}_{k}{\bm{e}}-\eta_{k}\left(\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}})+\Xi_{k+1}{\bm{e}}\right).$ (3.24)
Moreover, Assumption 3.1(3) illustrates that
$\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}{\bm{e}}\right\|\leq\frac{1}{\sqrt{d}}\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$ (3.25)
Then combined with Assumption 3.1(2), it directly follows from Lemma 2.11 that the interpolated process of the sequence $\{{\bm{Z}}_{k}{\bm{e}}\}$ is a perturbed solution of the differential inclusion (3.4). This completes the proof. ∎
Based on Proposition 3.7 and Theorem 2.12, we establish the convergence theorem, where we prove that the sequence $\{{\bm{Z}}_{k}\}$ converges to the stable set of (3.4) for our framework (DSM).
Theorem 3.8.
Suppose Assumption 3.1 and Assumption 3.2 hold. Then for the sequence $\{{\bm{Z}}_{k}\}$ generated by our framework (DSM), any cluster point of $\{{\bm{Z}}_{k}\}$ lies in $\{{\bm{Z}}\in\mathbb{R}^{m\times d}:{\bm{Z}}={\bm{Z}}{\bm{1}}_{d}^{\top},~{}{\bm{Z}}\in\mathcal{A}\}$ and the sequence $\{\Psi({\bm{z}}_{i,k})\}$ converges for each $i\in[d]$.
Proof.
Proposition 3.7 illustrates that the interpolated process of $\{{\bm{Z}}_{k}{\bm{e}}\}$ is a perturbed solution to the differential inclusion (3.4). Together with Assumption 3.2 and Theorem 2.12, we can conclude that any cluster point of $\{{\bm{Z}}_{k}{\bm{e}}\}$ lies in $\mathcal{A}$, and the sequence of function values $\{\Psi({\bm{Z}}_{k}{\bm{e}})\}$ converges. Moreover, notice that Proposition 3.5 illustrates that
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{e}}{\bm{1}}_{d}^{\top}\right\|=0.$ (3.26)
Therefore, we can conclude that any cluster point of $\{{\bm{Z}}_{k}\}$ lies in $\{{\bm{Z}}\in\mathbb{R}^{m\times d}:{\bm{Z}}={\bm{Z}}{\bm{1}}_{d}^{\top},~{}{\bm{Z}}\in\mathcal{A}\}$, and the sequence of function values $\{\Psi({\bm{z}}_{i,k})\}$ converges for each $i\in[d]$. ∎
4 Application: Decentralized Stochastic Subgradient Methods
In this section, we aim to illustrate the great flexibility of our proposed framework (DSM). We show that our proposed framework (DSM) encloses a wide range of popular decentralized stochastic subgradient methods, including DSGD, DSGD with momentum (DSGDm), and DSGD with gradient-tracking technique (DSGD-T). More importantly, we demonstrate that our theoretical results in Theorem 3.8 provide theoretical convergence guarantees for these decentralized stochastic subgradient methods, especially in the minimization of nonsmooth path-differentiable functions. Throughout this section, we make the following assumptions on (DOP).
Assumption 4.1.
  1. 1.
    For each $i\in[d]$, there exist functions $\{f_{i,j}:j\in[N_{i}]\}$ such that for any $i\in[d]$, it holds that
    $f_{i}({\bm{x}})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}f_{i,j}({\bm{x}}).$
    Moreover, for each $i\in[d]$ and $j\in[N_{i}]$, $f_{i,j}$ is definable and admits $\mathcal{D}_{f_{i,j}}$ as its definable conservative field.
  2. 2.
    $f_{i}({\bm{x}})$ is bounded from below.
Remark 4.2.
It is worth mentioning that Assumption 4.1 is mild and reasonable in practical scenarios. As demonstrated in Section 2, the class of definable functions is wide enough to enclose a wide range of the loss functions of common nonsmooth neural networks. Furthermore, [8] demonstrates that for the loss function of any neural network constructed by definable blocks, the results yielded by Automatic Differentiation (AD) algorithms are contained within a definable conservative field for the definable loss function. Moreover, for the objective functions $\{f_{i}\}$, we can choose their definable conservative fields as
$\mathcal{D}_{f_{i}}:=\sum_{j=1}^{N_{i}}\mathcal{D}_{f_{i,j}}({\bm{x}}).$ (4.1)
In addition, recall that $f({\bm{x}})=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ as defined in (2.1), we can choose its definable conservative field $\mathcal{D}_{f}$ as
$\mathcal{D}_{f}({\bm{x}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}})\right).$ (4.2)
Then by Proposition 2.5, we can conclude that the set $\{{f({\bm{x}})}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is a finite subset of $\mathbb{R}$.
4.1 Decentralized stochastic subgradient descent
In this subsection, we establish the global convergence for decentralized stochastic subgradient descent (DSGD) by showing that it fits into our proposed framework (DSM). The detailed algorithm of DSGD is presented in Algorithm 1. Intuitively, for each $i\in[d]$, the consensus term $\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}$ enforces ${\bm{x}}_{i,k}$ towards the mean over all the agents (i.e., $\frac{1}{d}\sum_{i\in[d]}{\bm{x}}_{i,k}$). Moreover, as ${\bm{d}}_{i,k}$ refers to the stochastic subgradient of $f_{i}$ at ${\bm{x}}_{i,k}$, the term $-\eta_{k}{\bm{d}}_{i,k}$ can be viewed as a single SGD step to minimize $f_{i}$ in the $i$-th agent.
Algorithm 1 DSGD for solving (DOP).
1:Initial point ${\bm{x}}_{0}\in\mathbb{R}^{n}$ and a mixing matrix ${\bm{W}}$.
2:for all $i\in[d]$ in parallel do
3:     Set $k\leftarrow 0$. Initialize ${\bm{x}}_{i,k}={\bm{x}}_{0}$
4:     while not terminated do
5:         Sample the index $j\in[N_{i}]$ uniformly at random;
6:         Compute ${\bm{d}}_{i,k}\in\mathcal{D}_{f_{i,j}}({\bm{x}}_{i,k})$;
7:         Choose the step-size $\eta_{k}$;
8:         Communicate and update the local variables ${\bm{x}}_{i,k}=\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{d}}_{i,k}$;
9:         Set $k\leftarrow k+1$;
10:     end while
11:end for
12:return ${\bm{X}}_{k}:=[{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}]$;
To establish the convergence properties of Algorithm 1, we first make the following assumptions based on Assumption 3.1.
Assumption 4.3.
  1. 1.
    ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix.
  2. 2.
    There exists a constant $M_{x}>0$ such that $\sup_{k\geq 0}\left\|{\bm{X}}_{k}\right\|\leq M_{x}$.
  3. 3.
    The sequence of step-sizes $\{\eta_{k}\}$ satisfies
    $\lim_{k\to+\infty}\eta_{k}\log(k)=0,\quad\sum_{i=0}^{+\infty}\eta_{k}=+\infty.$ (4.3)
Assumption 4.3(1) is identical to Assumption 3.1(1), while Assumption 4.3(2) assumes the uniform boundedness of the sequence $\{{\bm{X}}_{k}\}=\{[\bm{x}_{1,k},...,\bm{x}_{d,k}]\}$. Moreover, Assumption 4.3(3) illustrates that the step-sizes diminish in the rate of $o(1/\log(k))$, hence enabling great flexibility in choosing the step-sizes $\{\eta_{k}\}$ in practical scenarios. In addition, as demonstrated in [2, Proposition 1.3], Assumption 4.3(3) implies Assumption 3.1(3) almost surely. In the following theorem, we establish the global convergence of Algorithm 1 by showing that Algorithm 1 fits into our proposed framework with $\mathcal{H}=\mathcal{D}_{f}$, $\Psi=f$ and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
Theorem 4.4.
Suppose Assumption 4.1 and 4.3 hold for Algorithm 1. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-critical point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
Proof.
As demonstrated in Line 7 of Algorithm 1, the sequence $\{{\bm{X}}_{k}\}$ follows the update scheme,
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{D}}_{k},\qquad\text{ where }\quad{\bm{D}}_{k}=[{\bm{d}}_{1,k},\ldots,{\bm{d}}_{d,k}].$ (4.4)
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. Then we set ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{D}}_{k}-\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$. Then it is easy to verify that the update scheme (4.4) is equivalent to the update scheme (DSM). Now we verify the validity of Assumption 3.1 and Assumption 3.2. Assumption 3.1(1)-(2) directly follows from Assumption 4.3(1)-(2) and local boundedness of $\mathcal{D}_{f_{i}}$ in (4.1). Moreover, Assumption 3.1(3) follows from Assumption 4.3(3) and [2, Proposition 1.3]. In addition, from the definition of $\mathcal{D}_{f_{i}}$, we can conclude that
${\bm{H}}_{k}{\bm{e}}\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k})\subseteq\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k})\right).$ (4.5)
Therefore, by choosing $\mathcal{H}_{i}=\mathcal{D}_{f_{i}}$, we verify the validity of Assumption 3.1(4). Furthermore, according to [8, Section 6] and the path-differentiability of $f$, it follows that $f$ is the Lyapunov function of the differential inclusion
$\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}}),$ (4.6)
and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, Remark 4.2 illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.2 with $\mathcal{H}=\mathcal{D}_{f}$, $\Psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. Finally, from Theorem 3.8, we can conclude that almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and the sequence of function values $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎
4.2 Decentralized generalized SGD with momentum
Momentum is a technique to accelerate the subgradient descent method in the relevant direction and dampen oscillations. In the single-agent setting, [46] investigates the global convergence properties of generalized SGD with momentum, where an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is introduced to normalize the update directions of the primal variables. However, to the best of our knowledge, there is no existing work on developing its decentralized version. In this subsection, motivated by the generalized SGD with momentum analyzed in [46], we first introduce the decentralized generalized SGD with momentum (DGSGDm) and present the details in Algorithm 2. Different from DSGD in Algorithm 1, Algorithm 2 introduces the auxiliary variables $\{{\bm{y}}_{i,k}:k\in\mathbb{N}\}$ to track the first-order moment of the stochastic subgradients of $f_{i}$. Therefore, Algorithm 2 requires two rounds of communication to perform local consensus to $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$ and $\{{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}\}$, respectively. Moreover, similar to the generalized SGD analyzed in [46], Algorithm 2 introduces an auxiliary function $\phi$ to normalize the update directions of its primal variables $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$.
Remark 4.5.
The auxiliary function $\phi$ in Algorithm 2 determines how the update directions of the variables $\{{\bm{x}}_{k}\}$ are regularized. Moreover, Algorithm 2 yields different variants of DSGD-type methods with different choices of the auxiliary function $\phi$. For example, Algorithm 2 corresponds to DSGDm when choosing $\phi({\bm{y}})=\frac{1}{2}\left\|{\bm{y}}\right\|^{2}$. Furthermore, when we select $\phi({\bm{y}})=\left\|{\bm{y}}\right\|_{1}$, Algorithm 2 employs the sign map to regularize the update directions of the variables $\{{\bm{x}}_{k}\}$. Under such selection of $\phi$, we can view this algorithm as a decentralized variant of SignSGD analyzed in [3], and denoted it as DSignSGD. Interested readers could refer to [46] for more details on the choices of the auxiliary function $\phi$.
Algorithm 2 DGSGDm for solving (DOP).
1:Initial point ${\bm{x}}_{0}=[{\bm{x}}_{1,0},\ldots,{\bm{x}}_{d,0}]$, initial descent direction ${\bm{y}}_{0}=[{\bm{y}}_{1,0},\ldots,{\bm{y}}_{d,0}]$, momentum parameter $\tau>0$, an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$, and a mixing matrix ${\bm{W}}$.
2:for all $i\in[d]$ in parallel do
3:     Set $k\leftarrow 0$;
4:     while not terminated do
5:         Choose the step-size $\eta_{k}$;
6:         Communicate and update local variable ${\bm{x}}_{i,k+1}\in\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}\partial\phi({\bm{y}}_{i,k})$;
7:         Sample the index $j\in[N_{i}]$ uniformly at random;
8:         Compute ${\bm{d}}_{i,k+1}\in\mathcal{D}_{f_{i,j}}({\bm{x}}_{i,k+1})$;
9:         Communicate and update local descent direction
${\bm{y}}_{i,k+1}=(1-\tau\eta_{k})\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{y}}_{j,k}+\tau\eta_{k}{\bm{d}}_{i,k+1};$
10:         Set $k\leftarrow k+1$;
11:     end while
12:end for
13:return ${\bm{X}}_{k}:=[{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}]$;
Let ${\bm{Y}}_{k}=[{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}]$ at $k$-th iteration in Algorithm 2, we present the following theorem to establish the global convergence of Algorithm 2.
Theorem 4.6.
Suppose Assumption 4.1 and 4.3 hold for Algorithm 2. Moreover, we assume that the auxiliary function $\phi$ is convex and admits a unique minimizer at ${\bm{0}}$. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
Proof.
From Line 5 and Line 8 of Algorithm 2, we can conclude that the sequences $\{{\bm{X}}_{k}\}$ and $\{{\bm{Y}}_{k}\}$ follow the update scheme,
${\bm{X}}_{k+1}\in{}$ ${\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{T}}_{k},\quad{\bm{T}}_{k}\in[\partial\phi({\bm{y}}_{1,k}),\ldots,\partial\phi({\bm{y}}_{d,k})],$ (4.7)
${\bm{D}}_{k+1}={}$ $[{\bm{d}}_{1,k+1},\ldots,{\bm{d}}_{d,k+1}],$
${\bm{Y}}_{k+1}={}$ $(1-\tau\eta_{k}){\bm{Y}}_{k}{\bm{W}}+\tau\eta_{k}{\bm{D}}_{k+1}.$
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. Then by setting ${\bm{Z}}_{k}=\left[\begin{smallmatrix}{\bm{X}}_{k}\\ {\bm{Y}}_{k}\end{smallmatrix}\right]$, ${\bm{H}}_{k}=\mathbb{E}\left[\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]|\mathcal{F}_{k}\right]$, and $\Xi_{k+1}=\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]-{\bm{H}}_{k}$, we can conclude that the update scheme (4.7) can be reshaped as
${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}).$ (4.8)
Now we verify the validity of Assumption 3.1 and Assumption 3.2. Assumption 3.1(1) is identical to Assumption 4.3(1). Moreover, from the uniform boundedness of $\{{\bm{X}}_{k}\}$ in Assumption 4.3(1) and the local boundedness of $\{\mathcal{D}_{f_{i}}\}$, we can conclude that $\{{\bm{D}}_{k}\}$ is uniformly bounded almost surely. Together with the update scheme of ${\bm{Y}}_{k}$, we can conclude that $\left\|{\bm{Y}}_{k+1}\right\|\leq(1-\tau\eta_{k})\left\|{\bm{Y}}_{k}\right\|+\tau\eta_{k}\left\|{\bm{D}}_{k+1}\right\|$, hence there exists a constant $M_{Y}>0$ such that $\sup_{k\geq 0}\left\|{\bm{Y}}_{k}\right\|\leq\max\{\left\|{\bm{Y}}_{0}\right\|,\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\}<M_{Y}$. Then combining with local boundedness of $\partial\phi$, it can be referred that $\{{\bm{T}}_{k}\}$ is uniformly bounded almost surely. This illustrates that Assumption 3.1(2) holds. Moreover, by the definition of $\{\Xi_{k}\}$, it is straightforward to see that $\{\Xi_{k}\}$ is uniformly bounded almost surely. Then Assumption 4.3(3) and [2, Remark 1.5] illustrates that Assumption 3.1(3) holds almost surely for Algorithm 2. In addition, from the definition of $\mathcal{D}_{f_{i}}$, let $\delta_{k}=\left\|{\bm{X}}_{k+1}-{\bm{X}}_{k}\right\|_{\mathrm{F}}$ and
$\mathcal{H}_{i}({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f_{i}}({\bm{x}})\end{smallmatrix}\right],$ (4.9)
we can conclude that
${\bm{H}}_{k}{\bm{e}}\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}^{\delta_{k}}_{i}({\bm{x}}_{i,k},{\bm{y}}_{i,k}).$ (4.10)
Followed by the same techniques in Lemma 3.4 and Proposition 3.5, we can prove that $\delta_{k}\to 0$. This verifies the validity of Assumption 3.1(4). Furthermore, from the definition of $f$ in Assumption 4.1(1), we attain that
$\mathcal{H}({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f}({\bm{x}})\end{smallmatrix}\right].$ (4.11)
Then together with [46, Proposition 4.5] and the path-differentiability of $f$, we can conclude that $\Psi({\bm{x}},{\bm{y}})=f({\bm{x}})+\frac{1}{\tau}\phi({\bm{y}})$ is the Lyapunov function of the differential inclusion
$\left(\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t},\frac{\mathrm{d}{\bm{y}}}{\mathrm{d}t}\right)\in-\mathcal{H}({\bm{x}},{\bm{y}}),$ (4.12)
and admits $\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ as its stable set. Together with the definability of $f$ and $\mathcal{D}_{f}$, we can conclude that the set $\{\Psi({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ is finite in $\mathbb{R}$. Therefore, Assumption 3.2 holds for Algorithm 2. Finally, from Theorem 3.8, we can conclude that almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎
4.3 Decentralized SGD with gradient-tracking technique
The decentralized SGD method with gradient-tracking technique (DSGD-T) is another widely employed stochastic (sub)gradient method in decentralized learning. Algorithm 3 exhibits the detailed implementations of DSGD-T in minimizing nonsmooth path-differentiable functions. Different from the DSGD in Algorithm 1, DSGD-T introduces a sequence of auxiliary variables to track the average of local subgradients over all the agents. To the best of our knowledge, all the existing works focus on analyzing the convergence of DSGD-T with Clarke regular objective functions, and the convergence of DSGD-T for nonsmooth nonconvex functions remains unexplored. In this subsection, we prove that DSGD-T can be enclosed by our proposed framework (DSM). Therefore, based on our theoretical results exhibited in Theorem 3.8, we can establish the global convergence of DSGD-T in the minimization of nonsmooth path-differentiable functions.
Algorithm 3 DSGD-T for solving (DOP)
1:Initial point ${\bm{x}}_{0}=[{\bm{x}}_{1,0},\ldots,{\bm{x}}_{d,0}]$, initial descent direction ${\bm{v}}_{0}=[{\bm{v}}_{1,0},\ldots,{\bm{v}}_{d,0}]$ satisfying ${\bm{v}}_{0}{\bm{1}}_{d}=\sum_{i=1}^{d}{\bm{d}}_{0,i}$, where ${\bm{d}}_{i,0}\in\mathcal{D}_{f_{i,j_{i,0}}}({\bm{x}}_{i,0}),$ $j_{i,0}$ is a uniformly random index in $[N_{i}]$, and a mixing matrix ${\bm{W}}$.
2:for all $i\in[d]$ in parallel do
3:     Set $k\leftarrow 0$;
4:     while not terminated do
5:         Choose the step-size $\eta_{k}$;
6:         Communicate and update local variable ${\bm{x}}_{i,k+1}=\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{v}}_{i,k}$;
7:         Sample the index $j\in[N_{i}]$ uniformly at random;
8:         Compute ${\bm{d}}_{i,k+1}\in\mathcal{D}_{f_{i,j}}({\bm{x}}_{i,k+1})$;
9:         Communicate and update local descent direction
${\bm{v}}_{i,k+1}=\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j)({\bm{v}}_{j,k}+{\bm{d}}_{j,k+1}-{\bm{d}}_{j,k});$
10:         Set $k\leftarrow k+1$;
11:     end while
12:end for
13:return ${\bm{X}}_{k}:=[{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}]$;
In the rest of this subsection, we denote ${\bm{V}}_{k}=[{\bm{v}}_{1,k},\ldots,{\bm{v}}_{d,k}]$ at $k$-th iteration in Algorithm 3. Assumption 4.3 and 4.1 are also adopted for DSGD-T. In the following, we condense the assertion that DSGD-T can fit in our framework (DSM), along with convergence analysis, into a single theorem.
Theorem 4.7.
Suppose Assumption 4.1 and 4.3 hold for Algorithm 3. Then the iterative formula of $\{{\bm{X}}_{k}\}$ is a specific form of our proposed general framework (DSM). Moreover, for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
Proof.
From Line 5 - Line 8 of Algorithm 3, the sequences $\{{\bm{X}}_{k}\}$ and $\{{\bm{V}}_{k}\}$ are updated by the following scheme,
${\bm{X}}_{k+1}={}$ ${\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{V}}_{k},\quad{\bm{D}}_{k+1}={}[{\bm{d}}_{1,k+1},\ldots,{\bm{d}}_{d,k+1}],$ (4.13)
${\bm{V}}_{k+1}={}$ $({\bm{V}}_{k}+{\bm{D}}_{k+1}-{\bm{D}}_{k}){\bm{W}}.$
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. By setting ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{V}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{V}}_{k}-{\bm{H}}_{k}$, we can conclude that (4.13) coincides with the framework (DSM). Then we aim to check the validity of Assumption 3.1. From the uniform boundedness of $\{{\bm{x}}_{k}\}$ in Assumption 4.3(2), we can conclude that there exists $M_{D}>0$ such that $\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\leq M_{D}$ almost surely. Then by the update scheme of $\{{\bm{V}}_{k}\}$ in (4.13), it holds that
${\bm{V}}_{k+1}={\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}.$ (4.14)
Then from Proposition 2.2, there exists a constant $\alpha\in(0,1)$, such that $\left\|({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k}\right\|\leq 2\alpha^{k}$ holds for any $k\geq 0$. Therefore, almost surely, it holds that
$\sup_{k\geq 0}\left\|{\bm{V}}_{k}\right\|\leq\sup_{k\geq 0}\left\|{\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}\right\|$ (4.15)
[…]
Therefore, we can conclude that $\{{\bm{V}}_{k}\}$ is uniformly bounded almost surely, hence verifying the validity of Assumption 3.1(2). Moreover, the uniform boundedness of $\{\Xi_{k}\}$ and [2, Remark 1.5] illustrates that Assumption 3.1(3) holds almost surely for Algorithm 3. In addition, it is easy to verify that
${\bm{H}}_{k}{\bm{e}}=\mathbb{E}[{\bm{V}}_{k}{\bm{e}}|\mathcal{F}_{k}]=\mathbb{E}[{\bm{D}}_{k}{\bm{e}}|\mathcal{F}_{k}]\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k}),\quad\forall k\in\mathbb{N}.$ (4.16)
Therefore, by choosing $\mathcal{H}_{i}=\mathcal{D}_{f_{i}}$ for all $i\in[d]$, we verify the validity of Assumption 3.1(4). Furthermore, according to [8, Section 6] and Remark 4.2, we can assert that $f$ is the Lyapunov function of the differential inclusion
$\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}}),$ (4.17)
and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, the definability of $f$ and $\mathcal{D}_{f}$ illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.2 with $\mathcal{H}=\mathcal{D}_{f}$, $\Psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. Finally, from Theorem 3.8, we can conclude that almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎
5 Numerical Experiments
In this section, we evaluate the numerical performance of our analyzed decentralized stochastic subgradient methods in Section 4 for decentralized training nonsmooth neural networks. All numerical experiments in this section are conducted on a workstation equipped with two Intel(R) Xeon(R) Processors Gold 5317 CPUs (at 3.00GHz, 18M Cache) and 8 NVIDIA GeForce RTX 4090 GPUs under Ubuntu 20.04.1. We implement all decentralized algorithms with Python 3.8 and PyTorch 1.13.0 using NCCL 2.14.3 (CUDA 11.7) as the communication backend.
5.1 Testing problems and default setting
We implement image classification tasks by the decentralized training of the ResNet-50 [20] network on the CIFAR-10 and CIFAR-100 datasets [25], which both consist of $50000$ training samples and $10000$ test samples. The decentralized network topology is set as ring-structure with $8$ agents, and the mixing matrix is chosen as the Metropolis weight matrix [44]. In fact, in each experiment, we divide the training dataset into 8 equal parts, with each part serving as the local dataset for each agent. During any epoch of training, the local dataset of each agent itself remains unchanged, but the order of batches is reshuffled to ensure compliance with the format of our problem (DOP) and the randomness requirements of all algorithms in Section 4. Moreover, we set the number of epochs as $200$ and choose the batch size as $128$. The strategy for selecting step-size is to decrease it at the $60$-th, $120$-th and $160$-th epochs, with a decaying factor of $0.2$.
5.2 Performance comparison between DSGD, DSGDm, and DSignSGD
We first evaluate the numerical performance of DSGD (Algorithm 1), DSGDm (Algorithm 2 with $\phi=\frac{1}{2}\|{\bm{y}}\|^{2}$), and DSignSGD (Algorithm 2 with $\phi=\left\|{\bm{y}}\right\|_{1}$). All compared methods are executed five times with varying random seeds. In each test instance, all these compared methods are initialized with the same randomly generated initial points. In Algorithm 2, we choose the momentum parameter $\tau$ as $\frac{0.1}{\eta_{0}}$, and the initial descent direction is taken as ${\bm{Y}}_{0}=(1-\tau\eta_{0}){\bm{D}}_{0}{\bm{W}}+\tau\eta_{0}{\bm{D}}_{0}$. Additionally, we set the initial step-size $\eta_{0}$ of DSGD and DSGDm to be $0.2$, and the counterpart of DSignSGD to be $10^{-4}$. We present the numerical results of 3 decentralized subgradient methods in Figure 1, including the test accuracy, train loss, and the consensus error. It is worth mentioning that the consensus error at the $k$-th iteration is measured by $\frac{1}{\sqrt{d}}\|{\bm{X}}_{k}({\bm{I}}_{d}-{\bm{P}})\|_{\mathrm{F}}$. As illustrated in Figure 1, all the compared decentralized stochastic subgradient methods are able to train the ResNet-50 network to a high accuracy. Moreover, we can conclude from Figure 1(a)(d) that DSignSGD achieves similar test accuracy as DSGD and DSGDm, while showing superior performance to DSGD and DSGDm in the aspect of train loss from Figure 1(b)(e). Furthermore, in all the compared methods, the consensus errors diminish to 00 throughout the iterations. These results illustrate that our framework is able to yield efficient stochastic subgradient methods for decentralized training tasks.
(a) Test Accuracy, CIFAR-100
(b) Train Loss, CIFAR-100
(c) Consensus Error, CIFAR-100
(d) Test Accuracy, CIFAR-10
(e) Train Loss, CIFAR-10
(f) Consensus Error, CIFAR-10
Figure 1: Numerical performance comparison of 3 decentralized subgradient methods in training ResNet50 on CIFAR-10 and CIFAR-100 datasets.
5.3 Performance comparison between DSGD and DSGD-T
In this subsection, we evaluate the numerical performance of DSGD (Algorithm 1) and DSGD-T (Algorithm 3). Similar to the settings in Section 5.2, we repeat the numerical experiments for $5$ times with varying random seeds for DSGD and DSGD-T. For each random seed, DSGD and DSGD-T are executed with the same initial point ${\bm{x}}_{0}$. The initial step-size is taken as $0.2$ for both algorithms. The numerical results are presented in Figure 2. It can be concluded that DSGD-T demonstrates slightly better performance than DSGD in all aspects, especially in train loss. Therefore, we can conclude that DSGD-T achieves high efficiency in decentralized training tasks while enjoying convergence guarantees from our framework (DSM).
(a) Test Accuracy, CIFAR-100
(b) Train Loss, CIFAR-100
(c) Consensus Error, CIFAR-100
(d) Test Accuracy, CIFAR-10
(e) Train Loss, CIFAR-10
(f) Consensus Error, CIFAR-10
Figure 2: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-10 and CIFAR-100 datasets.
6 Conclusion
Decentralized stochastic subgradient methods play an important role in the decentralized training of deep neural networks, where the widely employed nonsmooth building blocks result in non-Clarke regular loss functions. As nearly all of the existing works on decentralized optimization focus on the Clarke regular functions, it is important to provide theoretical guarantees for these decentralized stochastic subgradient methods in the minimization of nonsmooth nonconvex functions. In this paper, we propose a general framework (DSM) to establish the convergence properties of decentralized stochastic subgradient methods for solving (DOP). We show that the sequence of iterates generated by our proposed framework asymptotically approximates the trajectory of its corresponding continuous-time noiseless differential inclusion. Then based on the ODE approaches, we establish the global convergence of our proposed framework (DSM) under mild conditions. We demonstrate the high flexibility of our proposed framework (DSM), in the sense that a wide range of decentralized stochastic subgradient methods fit into our proposed framework (DSM). In particular, we show that the framework (DSM) encloses various popular decentralized stochastic subgradient methods, including DSGD, DSGD-T, DSGDm, and DSignSGD. Based on our theoretical results, we establish global convergence for these decentralized stochastic subgradient methods in the minimization of nonsmooth nonconvex functions, especially in the decentralized training of nonsmooth neural networks. Preliminary numerical experiments verify the validity of our theoretical results and exhibit the high efficiency of these decentralized stochastic subgradient methods enclosed in our framework (DSM). Very recently, apart from the ODE approaches, the closed-measure approaches [4, 9, 26] are developed for establishing the convergence properties of stochastic subgradient methods, with relaxed conditions on the step-sizes and evaluation noises. To the best of our knowledge, all the existing results focus on nonsmooth optimization within a single agent. Hence, how to employ these closed-measure approaches for analyzing decentralized stochastic subgradient methods is an important direction of our future research.
Acknowledgement
The work of Professor Xin Liu was supported in part by the National Key R&D Program of China (2023YFA1009300), the National Natural Science Foundation of China (12125108, 12226008, 11991021, 11991020, 12021001, 12288201), Key Research Program of Frontier Sciences, Chinese Academy of Sciences (ZDBS-LY-7022), and CAS-Croucher Funding Scheme for Joint Laboratories “CAS AMSS-PolyU Joint Laboratory of Applied Mathematics: Nonlinear Optimization Theory, Algorithms and Applications”.